what is the pressure 12.5m below the ocean? Use the sp. gr. = 1.03 for salt water.


1
Expert's answer
2021-03-10T17:15:00-0500

The gauge pressure at any depth is given by,

Pg = ρgh

The atmospheric pressure:

"P_0 = 1.013 \times 10^5 \;Pa"

The pressure at this depth will be:

"P = P_0 + P_g \\

= 1.013 \times 10^5 + 1.03 \times 10^3 \times 9.8 \times 12.5 \\

= 2.27 \times 10^5 \;Pa"